OpenStudy (anonymous):

ok! If I have a table with X & P(x), how would I find the expected value?? x| 23|25|26|31|34|38 P(X)| 0.16|.09|0.18|0.12|0.24|0.21

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

@campbell_st @Loujoelou

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

would you do: X * P(X) and do it for all the values and ADD THEM all together to get the expected value??

OpenStudy (anonymous):

HELP guys?

OpenStudy (campbell_st):

well my understanding of expected value from probability is sum x*P(x) so 23*0.16 + 25*0.09 + .....

OpenStudy (anonymous):

yea that's what I was asking. so is that how u do it?

OpenStudy (anonymous):

cuz that's how I got 30.47

OpenStudy (campbell_st):

as its shown above I haven't done the calculation, you are summing 6 terms and doing a quick approximation, your answer seems reasonable

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

also:

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (mathmate):

Conditional probability: A=first throw, odd B=second throw, 1 or 6 \( P(B|A)=P(B\cap A) / P(A) \) You can see from the contingency table that there are 6 outcomes for \( A \cap B \) and 18 outcomes for A, so the probability is (6/36)/(18/36)=1/3